7. A 1000-kg car rounds a curve on a flat road of radius 50 m at a speed of 15 m/s (54 km/h). Will the car follow the curve, or will it skid? Assume:

(a) the pavement is dry and the coefficient of static friction is
μs = 0.60.



(b) the pavement is icy and μs = 0.25. (2 mark)
1
Expert's answer
2020-04-06T09:22:15-0400

The Newton's second law says

"ma=F."

For the circular motion on the flat road

"m\frac{v^2}{R}=\mu_smg."

Hence, the car will follow the curve if

"\mu_s\geq\frac{v^2}{gR}=\frac{15^2}{9.8\times 50}=0.46"

(a)

"\mu_s=0.60\geq 0.46"

The car will follow the curve.

(b)

"\mu_s=0.25\leq 0.46"

The car will skid the curve.